OpenStudy (anonymous):

MATH HELP PLEASE STUDY GUIDE a geostationary satellite is positioned 35,800 km above earths surface. It takes 24 hours to complete one orbit. The radius of Earth is about 6400 KM. a. what distance does the satellite travel in 1 hour? 3 hours? 2.5 hours? 25 hours? b. after how many hours has the satellite traveled 200,000 KM?

jimthompson5910 (jim_thompson5910):

The radius of the orbit is 6400+35800 = 42200 km

jimthompson5910 (jim_thompson5910):

so the circumference of the orbit is C = 2*pi*r C = 2*3.14*42200 C = ??

OpenStudy (anonymous):

The satellite travels in a circle of radius (35800 + 6400) = 42200 km. The circumference of a circle with this radius is the distance traveled by the satellite in 24 hours. Since the circumference of a circle is 2π times its radius, the satellite travels 2π*42200 ≈ 265150 km. If it takes 24 hours to travel that distance, its speed is 265150/24 ≈ 11047.9 km/hr To get your distances for 1 hr, 3 hr, etc., simply multiply the time in hours by the speed: x = vt. To answer your second question, set 200000 = 11047.9t and solve for t. Your answer will be in hours.

OpenStudy (anonymous):

In one hour the satellite will have traveled 266 km. In 3 hours, it will have traveled 800 km. In 2.5 hours, it will have traveled 666 km. And finally in 25 hours it will have traveled 6667 km. - See more at: http://www.chacha.com/question/if-a-satellite-is-positioned-35,800-km-above-earths-surface.-it-takes-24-hours-to-complete-one-orbit.-the-radius-of-earth-is-about-6,400-km.-what-distance-doe-the-satelite-travel-in-1-hour,-3-hours,-2.5-hours,-and-25-hours#sthash.KZgVbj6U.dpuf

OpenStudy (anonymous):

here's a website that I found this answer on :P

OpenStudy (anonymous):

@alli14344 thank you so much omg . <3 you're amazing.

OpenStudy (anonymous):

lol your welcome :) and thankyou

OpenStudy (anonymous):

and also thank you @jim_thompson5910 for helping! i wish i could've given u a medal but i guess i can't give more than one.

jimthompson5910 (jim_thompson5910):

lol that's fine, alli14344 did a much more complete job explaining I wanted to have you answer some of it so we can work through it together, but this works too